# Simulating Meaning, Nevermore! Introducing ICR: A Semiotic-Hermeneutic Metric for Evaluating Meaning in LLM Text Summaries

### v1

@@ 1 Introduction
Language meaning is dependent on signs, interrelations, and context. Saussure (1916) described this relationship as a linguistic sign system [59]; this system always contains the combination of a signifier (word) and a signified (conceptual meaning), with its sign (overall meaning) generated based on how the signifier and signified relate within a system of signs. This argument is easily illustrated within the poem The Raven, written by Edgar Allan Poe in 1845 [54]. The word “Nevermore” is a signifier (word), yet the word is repeated throughout The Raven, and its signified (conceptual meaning) changes with each new stanza. In one section of the poem, “Nevermore” refers to a lover never getting to see a lost love again, and a few stanzas later, “Nevermore” is used to refer to a person struggling to find comfort, and then it later is used to refer to a person experiencing eternal despair. Within the poem, Nevermore’s meaning changes, rather than staying the same, making Nevermore an example of a polysemy, which is one signifier (word) linked to multiple signifieds (meaning), within a single text. By contrast, large language models (LLMs) are often evaluated using automated evaluation metrics, which often treat signifiers (words) as static, context-independent units, which assumes that a word carries a fixed signified (conceptual meaning) [9]. As a result, LLM evaluation metrics might not always capture the context-dependent shifts in signifieds, producing outcomes that risk failing to capture the nuanced, relational and fluid meaning of signs present in texts like The Raven. It is along these lines that we argue the use of automated metrics as computational proxies for meaning could result in metric scores that do not always reflect the intended meaning of language within a reference text.
@@ 3.3 Analytical Overview
The final step is to achieve ICR is to calculate the results. To do so, researchers must calculate the conceptual presence or absence of concepts in the human-baseline and LLM-generate outputs. This calculation can produce a quantifiable interpretive rating of meaning accuracy (i.e., how closely GenAI outputs reflect RTA outputs) and truthfulness (i.e., whether GenAI outputs preserve contextual relationships and produce supported or misleading interpretations). After the outputs are evaluated, an ICR score, ranging from 0-1 can be assigned that calculates the frequency of true positives, false positives, true negatives, and false negatives in outputs. Overall, this metric is designed to capture gradations of alignment by using the RTA output as the baseline of accuracy in which the GenAI output can be compared and measured against.
@@
  • True negatives (TN): Not applicable here, or counted if non-relevant concepts are correctly excluded
The resulting ICR score quantifies conceptual alignment and contextual truthfulness, with higher scores indicating closer adherence to the human baseline and fewer misleading interpretations (see Appendix 4 for a more detailed example).
@@ D.1 Inductive Content Analysis of GenAI Outputs
    Assigning ICR Score The AI outputs were evaluated for:
    • True Positives (TP): GenAI outputs accurately reflected RTA themes (e.g., Flexibility, Career Emphasis)
    • False Positives (FP): GenAI outputs introduced interpretations not present in RTA (e.g., “all participants enjoy career success”)
    • False Negatives (FN): RTA themes omitted by GenAI model (e.g., nuanced emotional tension, guilt)
    • True Negatives (TN): Absence of irrelevant concepts
    An ICR score was calculated (e.g., 0.62), reflecting partial alignment, indicating that the AI captured some meaningful patterns but missed or distorted other nuanced aspects of human experience.

### v2

@@ 1 Introduction
Language meaning is dependent on signs, interrelations, and context. Saussure (1916) described this relationship as a linguistic sign system [59]; this system always contains the combination of a signifier (word) and a signified (conceptual meaning), with its sign (overall meaning) generated based on how the signifier and signified relate within a system of signs. This argument is easily illustrated within the poem The Raven, written by Edgar Allan Poe in 1845 [54]. The word “Nevermore” is a signifier (word), yet the word is repeated throughout The Raven, and its signified (conceptual meaning) changes with each new stanza. In one section of the poem, “Nevermore” refers to a lover never getting to see a lost love again, and a few stanzas later, “Nevermore” is used to refer to a person struggling to find comfort, and then it later is used to refer to a person experiencing eternal despair. Within the poem, Nevermore’s meaning changes, rather than staying the same, making Nevermore an example of a polysemy, which is one signifier (word) linked to multiple signifieds (meaning), within a single text [55]. In contrast, large language models (LLMs) are often evaluated using automated evaluation metrics, which often treat signifiers (words) as static, context-independent units, which assumes that a word carries a fixed signified (conceptual meaning) [9]. As a result, LLM evaluation metrics might not always capture the context-dependent shifts in signifieds, producing outcomes that risk failing to capture the nuanced, relational and fluid meaning of signs present in texts like The Raven. It is along these lines that we argue the use of automated metrics as computational proxies for meaning could result in metric scores that do not always reflect the intended meaning of language within a reference text.
@@ 3.3 Analytical Overview
The final step in calculating the ICR is to quantify the results. Researchers should evaluate the conceptual presence or absence of concepts in the human baseline and LLM-generated outputs using the accuracy formula:
$\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{True Positives}+\text{False Positives}+\text{True Negatives}+\text{False Negatives}}$
This calculation produces a quantifiable interpretive rating of meaning accuracy (i.e., how closely GenAI outputs reflect RTA outputs) and truthfulness (i.e., whether GenAI outputs preserve contextual relationships and produce supported or misleading interpretations). After evaluation, an ICR score ranging from 0 to 1 can be assigned. Higher values indicate stronger overall alignment with the human baseline, fewer errors (both false positives and false negatives), and better preservation of relevant concepts. This metric captures gradations of alignment by using the RTA output as the reference against which GenAI outputs are compared and measured.
@@
  • True negatives (TN): Not applicable here; no true negatives identified
In this example, the resulting ICR score is 0.6, which indicates that 60% of the concepts identified in the GenAI output align with those present in the human baseline, while 40% represent unsupported, potentially misleading, or hallucinatory additions. Interpreted in this way, the ICR score provides a quantitative indicator of conceptual alignment and contextual truthfulness between human-generated and GenAI-generated interpretations. A more detailed illustration of this scoring process is provided in Appendix 4.
@@ D.1 Inductive Content Analysis of GenAI Outputs
    Calculating the ICR Score
    The GenAI outputs were evaluated by comparing the conceptual content of the generated interpretations with the themes identified through Reflexive Thematic Analysis (RTA). Conceptual alignment was assessed using the following classification categories:
    • True Positives (TP): Concepts present in the RTA baseline that were accurately reflected in the GenAI output (e.g., Flexibility, Career Emphasis).
    • False Positives (FP): Concepts introduced by the GenAI output that were not present in the RTA baseline (e.g., “all participants enjoy career success”).
    • False Negatives (FN): Concepts identified in the RTA baseline but omitted from the GenAI output (e.g., nuanced emotional tension or feelings of guilt).
    • True Negatives (TN): The absence of unrelated or irrelevant conceptual content in the GenAI output.
    Using these classifications, an Inductive Conceptual Rating (ICR) score was calculated using an accuracy-based metric. For example, in the current dataset, the GenAI output correctly captured 5 relevant concepts (TP = 5), introduced 1 unsupported concept (FP = 1), omitted 1 relevant concept (FN = 1), and had no additional irrelevant concepts (TN = 0). Using the accuracy formula:
    $ICR(\text{Accuracy})=\frac{TP+TN}{TP+FP+TN+FN}=\frac{5+0}{5+1+0+1}=\frac{5}{7}\approx 0.714$
    This ICR score of approximately 0.714 reflects moderate-to-strong overall conceptual alignment between the GenAI interpretation and the human baseline; this indicates that most themes were captured accurately, while a smaller portion of identified concepts were either missed or incorrectly introduced. Higher ICR scores represent stronger correspondence with the human-generated analysis and fewer errors in conceptual interpretation.